Question 11 Suppose that a price-searcher firm had consumers who were all identical to each other. The individual consumers demand function is given by: qp- 40 -3P. The firm decides to try a second-degree price discrimination scheme. The first 18 units will have a price of $7.33. After that, any units a consumer purchases will be only $2.33. The firm has a constant marginal cost of $1.33 per unit. Calculate the consumer surplus. Tries remaining:2 Points out of 8.33 Flag question (Do not include a $ sign in your response. Round to the nearest two decimal places if necessary.) Answer: Check
Answer #1

Qd = 40 - 3P

P = 13.33 - 0.33Qd

When price is $7.33 quantity sold is 18 units

Consumer Surplus at this price = 0.5*(13.33-7.33)*(18-0)

CS1 = 0.5*6*18 = $ 54

When price is $2.33 quantity is

Q = 40 - 3*2.33 = 40 - 7 = 33

CS2=0.5*(33-18)*(7.33-2.33) = $ 37.5

Total consumer surplus = 54 + 37.5 = $ 91.50

Please contact if having any query thank you.